# On stable assignments generated by choice functions of mixed type

### v1

@@ 5 Additional properties of rotations
To show (ii), we first notice that any stable $x\in{\cal S}$ can be obtained from $x^{\rm min}$ by a finite sequence of rotational shifts with some admissible weights (this can be seen, e.g., from Lemma 4.12). If in this sequence, one and the same rotation $\rho_{L}$ (or, better to say, the same strong component $L$ in current active graphs) occurs several times, then we can combine the shifts related to these occurrences into one shift along $\rho_{L}$, summing up the corresponding rotational weights. (Here we use the fact that if a shift along $\rho_{L}$ is not full, then $L$ continues to be a maximal strong component in the current active graph, cf. the proof of Proposition 4.9.) Thus, the initial sequence can be transformed into another one, going from $x^{\rm min}$ to $x$ as before, so that each rotation in it occur exactly once. Then the rotations with non-maximal weights are related to different maximal strong components in the last graph $\overrightarrow{\Gamma}(x)$), and therefore, the number of these is $O(|V|)$. Now the desired estimate follows from (i). $\Box$
Remark 2. In fact, a slightly more refined analysis shows that the term $O(|E|)$ figured (twice) in Lemma 5.2 can be replaced by $2(|E|)$ (we omit details here). A similar result (with the bound $2|E|$) for SAP was established in [4] (extending an earlier, well known, result of this sort for the stable marriage problem in [9]).
@@
Example 3. As is mentioned in Remark 1 from Sect. 4, there exist rotations with “big” values. An example of this sort is shown in Fig. 1. Here we demonstrate an active graph $\Gamma$ with vertex parts $F=\{f_{1},f_{2},f_{3}\}$ and $W=\{w_{1},w_{2},w_{3}\}$, the edges active for $F$ ($W$) are drawn by bold (resp. dotted) lines, and the numbers on edges point out values of rotation $\rho$ (this is a normalized aligned $Z$-circulation); they vary in the interval $[-8,7]$.
@@
This rotation appears in the following instance of SMP: $G=\Gamma$; the weak orders at vertices of $G$ are defined as:
@@
For arbitrarily chosen numbers $0<a<b$, the upper bound $b(e)$ for each edge $e$ is equal to $b$, and the quota $q(v)$ of each vertex $v$ is equal to $3a$. Let $x(e):=a$ for all edges $e$ of $G$. One can see that: the assignment $x$ is fully filled for all vertices ($F^{=}(x)=F$ and $W^{=}(x)=W$); for all $w_{j}\in W$, the critical ties $\pi^{\rm c}_{w_{j}}(x)$ consist just of the dotted edges and they coincide with the heads $H_{w_{j}}(x)$; for all $f_{i}\in F$, the reduced heads $\widetilde{H}_{f_{i}}(x)$ consist of the bold edges. As a result, $x$ is stable. All edges are non-saturated by $x$, and one can see that the directed active graph $\overrightarrow{\Gamma}(x)$ has a unique strong component $L$ involving all edges. Then $\rho_{L}=\rho$. Applying ((5.1)), we obtain that the maximum rotational weight $\tau_{L}$ is equal to the minimum of $a/8$ and $(b-a)/7$.
@@ 6 The poset of rotations
Definition 7. Let $T$ be a sequence $x_{0},x_{1},\ldots,x_{N}$ of stable assignments such that each $x_{i}$ is obtained from $x_{i-1}$ by a rotational shift, i.e., $x_{i}=x_{i-1}+\lambda_{i}\rho_{L_{i}}$ (where $\rho_{L_{i}}$ is the rotation generated by a element (maximal strong component) $L_{i}$ of the set ${\cal L}^{\rm rot}(x_{i-1})$, and $\lambda_{i}$ is an admissible weight: $\lambda_{i}\leq\tau_{L_{i}}(x_{i-1})$; see Definition 6. We say that $T$ is non-expensive is all rotations $\rho_{L_{i}}$ are different (as vectors in ${\mathbb{R}}^{E}$). In particular, $T$ is non-expensive if all shifts involved are full, i.e., each $\lambda_{i}$ is the maximal rotational weight for $L_{i}$ (at the moment it applies). This always happens (in a non-expensive case) when $x_{0}=x^{\rm min}$ and $x_{N}=x^{\rm max}$; in this case we call $T$ a (complete) route. The set of all routs is denoted by ${\cal T}^{0}$.
When $x_{0}=x^{\rm min}$ and $x=x_{N}$ is arbitrary (resp. $x=x_{0}$ is arbitrary and $x_{N}=x^{\rm max}$), we call $T$ a lower (resp. upper) semi-route. The sequence of lower (resp. upper) sub-routes for $x\in{\cal S}$ is denoted by ${\cal T}^{\rm low}(x)$ (resp. ${\cal T}^{\rm up}(x)$). The set of rotations used to obtain a route or semi-route $T$ is denoted as ${\cal R}(T)$, and the set of pairs $(\lambda,\rho)$ formed by rotations $\rho\in{\cal R}(T)$ and their corresponding weights related to $T$ is denoted by $\Omega(T)$.
An important fact is that all routes use the same set of rotations and the same set of their weights. In other words, a strong component $L$ determining the rotation vector $\rho_{L}$ which is used in one route should necessarily appear in any other one and, moreover, for each rotation, the value of shift along it is always the same. Initially the phenomenon of this sort was revealed for rotations in the stable marriage problem in [9]; subsequently, this was established for other stability problems, in particular, for SAP in [4].
Generalizing a proving method from [4], we show a similar property for SMP.
@@
The set of pairs $\Omega(T)$ is the same for all routes $T\in{\cal T}^{0}$ (going from $x^{\rm min}$ to $x^{\rm max}$).
Proof  Let us say that a stable assignment $x\in{\cal S}$ is good if the set $\Omega(T)$ is the same for any upper semi-routes $T\in{\cal T}^{\rm up}(x)$; otherwise $x$ is called bad. We have to show that $x^{\rm min}$ is good.
Suppose that this is not so, and consider a maximal bad assignment $x$, in the sense that any $x^{\prime}$ obtained from $x$ by a rotational shift of maximal admissible weight is already good. Since $x$ is bad, in the set ${\cal T}^{\rm up}(x)$, there are two semi-routes $T^{\prime}=(x,x^{\prime},\ldots,x^{\rm max})$ and $T^{\prime\prime}=(x,x^{\prime\prime},\ldots,x^{\rm max})$ such that $\Omega(T^{\prime})\neq\Omega(T^{\prime\prime})$.
Let $x^{\prime}$ and $x^{\prime\prime}$ (occurring in $T^{\prime},T^{\prime\prime}$) be obtained from $x$ by rotational shifts by $\lambda^{\prime}\rho_{L^{\prime}}$ and $\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$, respectively. These shifts involve different maximal strong components $L^{\prime}$ and $L^{\prime\prime}$ which are simultaneously contained in the active graph $\overrightarrow{\Gamma}(x)$. These components are disjoint, and therefore, the above rotational shifts commute. Then we can form in ${\cal T}(x)$ two semi-routes; namely, $\widetilde{T}^{\prime}$ starting with the triple $x,x^{\prime},x^{\prime}+\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$, and $\widetilde{T}^{\prime\prime}$ starting with the triple $x,x^{\prime\prime},x^{\prime\prime}+\lambda^{\prime}\rho_{L^{\prime}}$, which further continue by the same upper semi-route $\widetilde{T}$ starting with $x+\lambda^{\prime}\rho_{L^{\prime}}+\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$.
@@
Note that from this proposition we can easily obtain the following
Corollary 6.2
For any $x,y\in{\cal S}$ with $x\succ y$, the sets of pairs $(\lambda,\rho)$ used to form non-expensive sequences (“middle semi-routes”) from $x$ to $y$ are the same.
@@
Definition 8.  A function $\lambda:{\cal R}\to{\mathbb{R}}_{+}$ is called closed if $\lambda(\rho)\leq\tau(\rho)$ for all $\rho\in{\cal R}$, and, in addition, $\lambda(\rho)>0$ and $\rho^{\prime}\lessdot\rho$ imply $\lambda(\rho^{\prime})=\tau(\rho^{\prime})$. A closed function $\lambda$ is called fully closed if for each $\rho\in{\cal R}$, the value $\lambda(\rho)$ is either 0 or $\tau(\rho)$. The set of closed functions for $({\cal R},\tau,\lessdot)$ is denoted by $\mathfrak{C}$. It forms a distributive lattice in which for closed $\lambda,\lambda^{\prime}$, the meet $\lambda\wedge\lambda^{\prime}$ and join $\lambda\vee\lambda^{\prime}$ are naturally defined by taking the coordinate-wise minimum and maximum, respectively.
Since the maximal weights $\tau(\rho)$ for all $\rho\in{\cal R}$ are strictly positive, we observe that for a closed function $\lambda$, both sets $I^{+}(\lambda):=\{\rho\colon\lambda(\rho)>0\}$ and $I^{-}(\lambda):=\{\rho\colon\lambda(\rho)=\tau(\rho)\}$ form ideals of the poset $({\cal R},\lessdot)$ (i.e. $I=I^{+}(\lambda)$ is such that $\rho\in I$ and $\rho^{\prime}\lessdot\rho$ imply $\rho^{\prime}\in I$, and similarly for $I^{-}(\lambda)$).
A relationship between the stable assignments and closed functions is established as follows. For $x\in{\cal S}$, consider a lower semi-route $T=(x^{\rm min}=x_{0},x_{1},\ldots,x_{N}=x)$, and for $i=1,\ldots,N$, let $x_{i}$ be obtained from $x_{i-1}$ by a rotational shift $\lambda_{i}\rho_{i}$. By Corollary 6.2, the set $\Omega(T)=\{(\lambda_{i},\rho_{i})\colon i=1,\ldots,N)\}$ does not depend on a semi-route in ${\cal T}^{\rm low}(x)$. Moreover, the function $\lambda:{\cal R}\to{\mathbb{R}}_{+}$, where $\lambda(\rho)$ is defined to be $\lambda_{i}$ if $\rho=\rho_{i}$, $i=1,\ldots,N$, and 0 otherwise, is closed. (Indeed, if $\rho=\rho_{i}$ and $\rho^{\prime}\lessdot\rho$, then $\rho^{\prime}$ must be used in any semi-route in ${\cal T}^{\rm low}(x)$, and its generating strong component $L^{\prime}$ cannot be in ${\cal L}^{\rm rot}(x_{i})$ (for otherwise $\rho^{\prime}$ and $\rho$ can be applied in any order). It follows that $\lambda(\rho^{\prime})=\tau(\lambda^{\prime})$, as required.) We define this closed function by $\omega(x)$.
@@
To show this, consider a middle (non-expensive) semi-route $X=(z=x_{0},x_{1},\ldots,x_{N})$ from $z:=x\curlywedge y$ to $x$, and a middle semi-route $Y=(z=y_{0},y_{1},\ldots,y_{M})$ from $z$ to $y$. Let ${\cal R}(X)=(\rho_{1},\ldots,\rho_{N})$ and ${\cal R}(Y)=(\rho^{\prime}_{1},\ldots,\rho^{\prime}_{M})$ be the corresponding sequences of rotations, and let $\lambda_{i}$ ($\lambda^{\prime}_{j}$) be the corresponding weight of $\rho_{i}$ (resp. $\rho^{\prime}_{j}$) used in $X$ (resp. $Y$). (By Corollary 6.2, the set of rotations $\rho_{i}$ in ${\cal R}(X)$, along with their weights $\lambda_{i}$, depend only on the end assignments $z$ and $x$. And similarly for ${\cal R}(Y)$.)
One can realize that the first relation in (6.1) follows from the property that
@@
If $i=j=1$ (i.e., $\rho$ is the first rotation in both ${\cal R}(X)$ and ${\cal R}(Y)$), then the assignment $z^{\prime}:=z+\min\{\lambda_{1},\lambda^{\prime}_{1}\}\,\rho$ satisfies $z^{\prime}\succ z$, $z^{\prime}\preceq x$ and $z^{\prime}\preceq y$, contrary to the fact that $z$ is the maximal lower bound for $x,y$. So we may assume that $i>1$ and $j\geq 1$.
@@
Claim. Let $\widetilde{x}\in{\cal S}$ and let $\rho$ and $\rho^{\prime}$ be rotations such that $\rho$ is applicable to $\widetilde{x}$, and $\rho^{\prime}$ is applicable to $\widetilde{x}^{\prime}:=\widetilde{x}+\lambda\rho$ for an admissible weight $\lambda$ for $\rho$. Then exactly one of the following is valid: (a) $\rho^{\prime}$ is applicable to $\widetilde{x}$, or (b) $\rho\lessdot\rho^{\prime}$.
This Claim will be proved later. Assuming validity of it, we argue as follows. Let $k$ be maximum subject to $k<i$ and $\rho_{k}\approx\rho_{i}$. First suppose that $k=i-1$. Then, by (a) in the Claim applied to $\widetilde{x}=x_{i-2}$, the rotations $\rho^{\prime}=\rho_{i-1}$ and $\rho=\rho_{i}$ commute, and we can update $X$ by swapping $\rho^{\prime}$ and $\rho$. This gives the semi-route $X^{\prime}=(x_{0},\ldots,x_{i-2},x^{\prime},x_{i},\ldots,x_{N})$, where $x^{\prime}:=x_{i-2}+\lambda_{i}\rho_{i}$, and in the corresponding sequence ${\cal R}(X^{\prime})$ the rotation $\rho$ is disposed earlier than in ${\cal R}(X)$. We can continue to further “lower” $\rho$ if $\rho_{i-2}\approx\rho$, and so on.

### v2

@@ 5 Additional properties of rotations
It can be seen (e.g., from Lemma 4.14) that $x^{\rm max}$ can be obtained from any $x\in{\cal S}$ by a finite sequence of rotational shifts with some admissible weights. If in this sequence, one and the same rotation $\rho_{L}$ (equivalently, the same maximal strong component $L$ in current active graphs) occurs several times, then we can combine the shifts related to these occurrences into one shift along $\rho_{L}$, summing up the rotational weights. (Here we use the fact that if a shift along $\rho_{L}$ is not full, then $L$ continues to be a maximal strong component in the current active graph, cf. the proof of Proposition 4.11.) Thus, the initial sequence can be transformed into another one, going from $x$ to $x^{\rm max}$ as well, so that each rotation in it occur exactly once. Then it is easy to see that all rotational shifts are full, and now the result follows from (i). $\Box$
Remark 2. Note that in general the set ${\cal S}$ of stable assignments in SMP need not be convex in ${\mathbb{R}}^{E}$. (This is so already for SAP with $c\equiv 1$ and $q\equiv 1$. For example, take as $G$ the 6-cycle $(v_{0},e_{1},v_{1},\ldots,e_{6},v_{6}=v_{0})$ and add to it edge $a=v_{1}v_{4}$. Assign $e_{i}<e_{i+1}$ for $i=1,\ldots,6$ (letting $e_{7}:=e_{1}$), $e_{1}<a<e_{2}$ and $e_{4}<a<e_{5}$. Then the incidence vectors of $\{e_{1},e_{3},e_{5}\}$ and $\{e_{2},e_{4},e_{6}\}$ give stable allocations but their half-sum $x$ does not, since the edge $a$ is blocking for $x$.) However, using Proposition 4.11, one can give the following geometric description of this set via rotations. Let ${\cal S}^{\rm full}$ be the set of stable assignments that can be obtained from $x^{\rm min}$ by a sequence of full rotational shifts. Then ${\cal S}$ is the union, over $x\in{\cal S}^{\rm full}$, of the rectangular parallelepipeds $x+Q(x)$, where $Q(x)$ is the Minkowsky sum of orthogonal segments $[0,\tau_{L}\rho_{L}]$, $L\in{\cal L}^{\rm rot}(x)$.
@@
Example 3. As is mentioned in Remark 1 from Sect. 4, there exist rotations with “big” values. An example is illustrated in Fig. 1. Here we draw an active graph $\Gamma$ with vertex parts $F=\{f_{1},f_{2},f_{3}\}$ and $W=\{w_{1},w_{2},w_{3}\}$, the edges active for $F$ ($W$) are drawn by bold (resp. dotted) lines, and the numbers on edges point out values of rotation $\rho$ (this is a normalized aligned $Z$-circulation); they vary from $-8$ through 7.
@@
Note that this rotation appears in the following instance of SMP: $G=\Gamma$; the weak orders at vertices of $G$ are defined as:
@@
For two arbitrarily chosen numbers $0<a<b$, the upper bound $b(e)$ for each edge $e$ is equal to $b$, and the quota $q(v)$ of each vertex $v$ is equal to $3a$. Let $x(e):=a$ for all edges $e$ of $G$. One can see that: all vertices in $F^{=}(x)=F$ and $W^{=}(x)=W$ are fully filled by $x$; for all $w_{j}\in W$, the critical ties $\pi^{\rm c}_{w_{j}}(x)$ consist just of the dotted edges and they coincide with the heads $H_{w_{j}}(x)$; for all $f_{i}\in F$, the heads $D_{f_{i}}(x)$ consist of the bold edges. As a result, $x$ is stable. All edges are unsaturated by $x$, and one can see that the directed active graph $\overrightarrow{\Gamma}(x)$ has a unique strong component $L$ involving all edges. Then $\rho_{L}=\rho$. Applying ((5.1)), we obtain that the maximum rotational weight $\tau_{L}(x)$ is equal to the minimum of $a/8$ and $(b-a)/7$.
@@
Remark 3. We can extend the above example as follows. Take copies $\Gamma^{1},\ldots,\Gamma^{k}$ of the graph $\Gamma$, denoting the copies of vertices $f_{i}$ and $w_{i}$ ($i=1,2,3$) in the graph $\Gamma^{j}$ ($j=1,\ldots,k$) as $f_{i}^{j}$ and $w_{i}^{j}$, respectively. For $j=1,\ldots,k-1$, identify the vertex $w_{1}^{j}$ of $\Gamma^{j}$ with the vertex $w_{3}^{j+1}$ of $\Gamma^{j+1}$, denoting the obtained vertex as $\widehat{w}^{j}$. Let $\widehat{\Gamma}$ be the resulting graph, the union of $\Gamma^{1},\ldots,\Gamma^{k}$. (In particular, $\widehat{w}^{1}$ has three $F$-active (“bold”) edges, namely, $f_{1}^{1}w_{1}^{1},\,f_{3}^{1}w_{1}^{1},\,f^{2}_{2}w_{3}^{2}$, and three $W$-active (“dotted”) edges, namely, $f_{2}^{1}w_{1}^{1},\,f_{1}^{2}w_{3}^{2},\,f_{3}^{2}w_{3}^{2}$ (where $w_{1}^{1}=w_{3}^{2}=\widehat{w}^{1}$).) Let $\widehat{\rho}$ be the rotation generated by $\widehat{\Gamma}$, and $\rho^{j}$ its restriction on $\Gamma^{j}$. One can see that $\rho^{1}$ coincides with the rotation $\rho$ for $\Gamma$ (taking into account that $\rho^{1}$ is determined by the balance and align conditions on the five vertices $f_{1}^{1},f_{2}^{1},f_{3}^{1},w_{2}^{1},w_{3}^{1}$). Then $\widehat{\rho}(f_{2}^{1}\widehat{w}^{1})=\rho(f_{2}w_{1})=-8$, which implies that $\widehat{\rho}(f_{1}^{2}\widehat{w}^{1})=\widehat{\rho}(f_{3}^{2}\widehat{w}^{1})=\widehat{\rho}(f_{2}^{1}\widehat{w}^{1})=-8$. Since $\rho(f_{1}w_{3})=\rho(f_{3}w_{3})=-2$, we obtain that $\rho^{2}$ is expressed as $4\rho$ (under identifying $\Gamma^{2}$ with $\Gamma$). Arguing similarly for $j=2$, we conclude that $\rho^{3}$ is expressed as $4\rho^{2}$, or as $16\rho$. Eventually, $\rho^{k}$ is expressed as $4^{k-1}\rho$.
Thus, the rotation $\widehat{\rho}$ in our construction has the $\ell_{\infty}$-norm $2\cdot 4^{k-1}$, whereas the graph $\widehat{\Gamma}$ has $15k+$ 1 vertices, thus providing an exponential grow of the rotation norms.
@@ 6 The poset of rotations
Definition 7. Let $T$ be a sequence $x_{0},x_{1},\ldots,x_{N}$ of stable assignments such that each $x_{i}$ is obtained from $x_{i-1}$ by a rotational shift, i.e., $x_{i}=x_{i-1}+\lambda_{i}\rho_{L_{i}}$ (where $\rho_{L_{i}}$ is the rotation generated by an element (maximal strong component) $L_{i}$ of the set ${\cal L}^{\rm rot}(x_{i-1})$, and $\lambda_{i}$ is an admissible weight: $\lambda_{i}\leq\tau_{L_{i}}(x_{i-1})$; see Definition 6. We say that $T$ is non-expensive is all rotations $\rho_{L_{i}}$ are different (as vectors in ${\mathbb{R}}^{E}$). In particular, $T$ is non-expensive if all involved shifts are full, in the sense that each $\lambda_{i}$ is the maximal rotational weight for $L_{i}$ at the moment it applies. A non-expensive $T$ with $x_{0}=x^{\rm min}$ and $x_{N}=x^{\rm max}$ is of most interest for us; such a $T$ is called a (complete) route. The set of all routs is denoted by ${\cal T}^{0}$.
In what follows we will deal with only non-expensive sequences $T=(x_{0},\ldots,x_{N})$. When $x_{0}=x^{\rm min}$ and $x=x_{N}$ is arbitrary (resp. $x=x_{0}$ is arbitrary and $x_{N}=x^{\rm max}$), we call $T$ a lower (resp. upper) semi-route. The set of lower (resp. upper) semi-routes for $x\in{\cal S}$ is denoted by ${\cal T}^{\rm low}(x)$ (resp. ${\cal T}^{\rm up}(x)$). The set of rotations used to form a route or semi-route $T$ is denoted as ${\cal R}(T)$. The set of pairs $(\lambda,\rho)$ formed by rotations $\rho\in{\cal R}(T)$ and their weights $\lambda$ used for $T$ is denoted by $\Omega(T)$.
An important fact is that all routes use the same set of rotations along with their weights. In other words, a strong component $L$ determining the rotation $\rho_{L}$ which is used in one route should necessarily appear in any other one and, moreover, each rotation uses the same weight in all occurrences. Initially the phenomenon of this sort was revealed for rotations in the stable marriage problem in [12]; subsequently, this was established for other stability problems, in particular, for SAP in [6].
Borrowing a method of proof for SAP from [6], we show a similar property for SMP.
@@
The set of pairs $\Omega(T)$ is the same for all routes $T\in{\cal T}^{0}$.
Proof  Let us say that a stable assignment $x\in{\cal S}$ is good if the set $\Omega(T)$ is the same for all upper semi-routes $T\in{\cal T}^{\rm up}(x)$; otherwise $x$ is called bad. We have to show that $x^{\rm min}$ is good.
Suppose this is not so, and consider a maximal bad assignment $x$, in the sense that any $x^{\prime}$ obtained from $x$ by a rotational shift of maximal admissible weight is already good. Since $x$ is bad, in the set ${\cal T}^{\rm up}(x)$, there are two semi-routes $T^{\prime}=(x,x^{\prime},\ldots,x^{\rm max})$ and $T^{\prime\prime}=(x,x^{\prime\prime},\ldots,x^{\rm max})$ such that $\Omega(T^{\prime})\neq\Omega(T^{\prime\prime})$.
Let $x^{\prime}$ and $x^{\prime\prime}$ (occurring in $T^{\prime},T^{\prime\prime}$) be obtained from $x$ by rotational shifts by $\lambda^{\prime}\rho_{L^{\prime}}$ and $\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$, respectively. These shifts involve different maximal strong components $L^{\prime}$ and $L^{\prime\prime}$ which are simultaneously contained in the active graph $\overrightarrow{\Gamma}(x)$. These components are disjoint, and the above rotational shifts commute. Then we can form in ${\cal T}^{\rm up}(x)$ two semi-routes; namely, $\widetilde{T}^{\prime}$ starting with the triple $x,x^{\prime},x^{\prime}+\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$, and $\widetilde{T}^{\prime\prime}$ starting with the triple $x,x^{\prime\prime},x^{\prime\prime}+\lambda^{\prime}\rho_{L^{\prime}}$, which further continue using the same upper semi-route $\widetilde{T}$ starting with $x+\lambda^{\prime}\rho_{L^{\prime}}+\lambda^{\prime\prime}\rho_{L^{\prime\prime}}$.
@@
The above proposition we can extended as follows (by arguing in a similar way).
Proposition 6.2
For any $x,y\in{\cal S}$ with $x\succ y$, the sets of pairs $(\lambda,\rho)$ used to form non-expensive sequences (“middle semi-routes”) from $x$ to $y$ are the same. $\Box$
@@
Definition 8.  A function $\lambda:{\cal R}\to{\mathbb{R}}_{+}$ satisfying $\lambda(\rho)\leq\tau(\rho)$ for all $\rho\in{\cal R}$ is called closed if $\lambda(\rho)>0$ and $\rho^{\prime}\lessdot\rho$ imply $\lambda(\rho^{\prime})=\tau(\rho^{\prime})$. A closed function $\lambda$ is called fully closed if for each $\rho\in{\cal R}$, the value $\lambda(\rho)$ is either 0 or $\tau(\rho)$. The set of closed functions for $({\cal R},\tau,\lessdot)$ is denoted by $\mathfrak{C}=\mathfrak{C}({\cal R},\tau,\lessdot)$. It forms a distributive lattice in which for closed $\lambda,\lambda^{\prime}$, the meet $\lambda\wedge\lambda^{\prime}$ and join $\lambda\vee\lambda^{\prime}$ are naturally defined by taking the coordinate-wise minimum and maximum, respectively.
Since the maximal weights $\tau(\rho)$ for all $\rho\in{\cal R}$ are strictly positive, we observe that for a closed function $\lambda$, both sets $I^{+}(\lambda):=\{\rho\colon\lambda(\rho)>0\}$ and $I^{-}(\lambda):=\{\rho\colon\lambda(\rho)=\tau(\rho)\}$ form ideals of the poset $({\cal R},\lessdot)$ (i.e. $\rho\in I^{+}(\lambda)$ and $\rho^{\prime}\lessdot\rho$ imply $\rho^{\prime}\in I^{+}(\lambda)$, and similarly for $I^{-}(\lambda)$).
A relationship between the stable assignments and closed functions is established as follows. For $x\in{\cal S}$, consider a lower semi-route $T=(x^{\rm min}=x_{0},x_{1},\ldots,x_{N}=x)$, and for $i=1,\ldots,N$, let $x_{i}$ be obtained from $x_{i-1}$ by a rotational shift $\lambda_{i}\rho_{i}$. By Proposition 6.2, the set $\Omega(T)=\{(\lambda_{i},\rho_{i})\colon i=1,\ldots,N)\}$ does not depend on a semi-route in ${\cal T}^{\rm low}(x)$. Moreover, the function $\lambda:{\cal R}\to{\mathbb{R}}_{+}$, where $\lambda(\rho)$ is defined to be $\lambda_{i}$ if $\rho=\rho_{i}$, $i=1,\ldots,N$, and 0 otherwise, is closed. (Indeed, if $\rho=\rho_{i}$ and $\rho^{\prime}\lessdot\rho$, then $\rho^{\prime}$ must be used in any semi-route in ${\cal T}^{\rm low}(x_{i})$, and its generating strong component $L^{\prime}$ cannot be in ${\cal L}^{\rm rot}(x_{i})$, for otherwise $\rho^{\prime}$ and $\rho$ can be applied in any order. It follows that $\lambda(\rho^{\prime})=\tau(\rho^{\prime})$, as required.) We define this closed function by $\omega(x)$.
@@
To show the left equality, consider a middle (non-expensive) semi-route $X=(z=x_{0},x_{1},\ldots,x_{N})$ from $z:=x\curlywedge y$ to $x$, and a middle semi-route $Y=(z=y_{0},y_{1},\ldots,y_{M})$ from $z$ to $y$. Let ${\cal R}(X)=(\rho_{1},\ldots,\rho_{N})$ and ${\cal R}(Y)=(\rho^{\prime}_{1},\ldots,\rho^{\prime}_{M})$ be the corresponding sequences of rotations, and let $\lambda_{i}$ ($\lambda^{\prime}_{j}$) be the corresponding weight of $\rho_{i}$ (resp. $\rho^{\prime}_{j}$) used in $X$ (resp. $Y$). (By Proposition 6.2, the set of rotations $\rho_{i}$ in ${\cal R}(X)$, along with their weights $\lambda_{i}$, depend only on the end assignments $z$ and $x$. And similarly for ${\cal R}(Y)$.)
One can realize that the left equality in (6.1) follows from the property that
@@
If $i=j=1$ (i.e., $\rho$ is the first rotation in both ${\cal R}(X)$ and ${\cal R}(Y)$), then $z^{\prime}:=z+\min\{\lambda_{1},\lambda^{\prime}_{1}\}\,\rho$ satisfies $z^{\prime}\succ z$, $z^{\prime}\preceq x$ and $z^{\prime}\preceq y$, contrary to the fact that $z$ is the maximal lower bound for $x,y$. So we may assume that $i>1$ and $j\geq 1$.
@@
Claim. Let $\widetilde{x}\in{\cal S}$ and let $\rho$ and $\rho^{\prime}$ be rotations such that $\rho$ is applicable to $\widetilde{x}$, and $\rho^{\prime}$ is applicable to $\widetilde{x}^{\prime}:=\widetilde{x}+\lambda\rho$ for an admissible weight $\lambda>0$ for $\rho$. Then exactly one of the following is valid: (a) $\rho^{\prime}$ is applicable to $\widetilde{x}$, or (b) $\rho\lessdot\rho^{\prime}$.
This Claim will be proved later. Assuming validity of it, we argue as follows. Let $k$ be maximum subject to $k<i$ and $\rho_{k}\approx\rho_{i}$. First suppose that $k=i-1$. Then, by (a) in the Claim applied to $\widetilde{x}=x_{i-2}$, the rotations $\rho^{\prime}=\rho_{i-1}$ and $\rho=\rho_{i}$ commute, and we can update $X$ by swapping $\rho^{\prime}$ and $\rho$. This gives the semi-route $X^{\prime}=(x_{0},\ldots,x_{i-2},x^{\prime},x_{i},\ldots,x_{N})$, where $x^{\prime}:=x_{i-2}+\lambda_{i}\rho_{i}$, and in the corresponding sequence ${\cal R}(X^{\prime})$ the rotation $\rho$ is disposed earlier than in ${\cal R}(X)$. We can continue to “lower” $\rho$ further if $\rho_{i-2}\approx\rho$, and so on.